Status Report

Hubble Space Telescope Daily Report #2824 8 March 2001

By SpaceRef Editor
March 8, 2001
Filed under ,



PERIOD COVERED: 0000Z (UTC) 03/07/01 – 0000Z (UTC) 03/08/01

Daily Status Report as of 067/0000Z


1.1 Completed WF/PC-2 8823 (Decontaminations and Associated
Observations Pt. 2/4)

The WF/PC-2 was used for the monthly WFPC2 decons. Also included
were instrument monitors tied to decons: photometric stability check, focus
monitor, pre- and post-decon internals {bias, intflats, kspots, & darks},
UV throughput check, VISFLAT sweep, and internal UV flat check. There were
no reported problems.

1.2 Completed Four Sets of WF/PC-2 8811 (Cycle 9 Standard Darks)

The WF/PC-2 was used to obtain dark frames every week in order to
provide data for the ongoing calibration of the CCD dark current rate, and
to monitor and characterize the evolution of hot pixels. Over an extended
period these data will also provide a monitor of radiation damage to the
CCDs. The proposal completed with no reported problems.

1.3 Completed Two Sets of STIS/CCD/MA2 8651 (UV Spectroscopy in the
Magellanic Bridge: A Typical QSO Absorption Line System?)

The Space Telescope Imaging Spectrograph (CCD and MA2) was used to
perform echelle UV absorption line spectroscopy through the Magellanic
Bridge {near the SMC} to probe the temperature, ionization, and kinematic
structure in a metal- poor environment like those observed in the early
universe. The sightline toward the QSO B0312-770 is especially interesting
because of the proximity of the absorber and the radio loud background
source. The observations completed nominally.

1.4 Completed Three Sets of WF/PC-2 8805 (POMS Test Proposal: WFII
Parallel Archive Proposal Continuation)

The WF/PC-2 was used to perform a generic target version of the
WFPC2 Archival Pure Parallel program. The program will be used to take
parallel images of random areas of the sky, following the recommendations
of the Parallels Working Group. The proposal completed with no reported

1.5 Completed Two Sets of WF/PC-2 8817 (Cycle 9 INTFLAT Sweeps and
Linearity Test)

The WF/PC-2 was used to obtain INTFLAT observations in order to
monitor the pixel to pixel flatfield response and provide a linearity
check. The INTFLAT sequences, to be done once during the year, are similar
to those from the Cycle 8 program 8448. The images will provide a backup
database in the event of complete failure of the VISFLAT lamp as well as
allow monitoring of the gain ratios. The sweep is a complete set of
internal flats, cycling through both shutter blades and both gains. The
linearity test consists of a series of INTFLATs in F555W, in each gain and
each shutter. There was no anomalous activity.

1.6 Completed STIS/CCD/MA2 8770 (The Nucleosynthesis of Boron –
Benchmarks for the Galactic Disk)

The Space Telescope Imaging Spectrograph (CCD and MA2) was used to
determine boron (B) abundance. In spite of several HST cycles of B
abundance determinations, we do not know the true, ambient, present-day
abundance of B. Only two stars {with undepleted Li and Be} have been
observed to derive the Population I true B abundance. Past cycles have
focused on {1} halo stars with low metallicity and {2} solar-type stars
with Li and/or Be deficiencies. In this proposal, the stellar abundances
of B will be determined, measured from the B I 2497 Angstrom line from STIS
echelle spectra, of main sequence F-G stars that have retained their full
initial abundances of boron. Our target stars are those for which
ground-based observations show that beryllium is undepleted. These new
abundances will map evolution of the B abundance in the Galactic disk in
the metallicity range from one-third solar up to solar. Then the B, and
previously determined Li and Be abundances, will be used to improve
understanding of the nucleosynthesis of B. There were no reported anomalies.


2.1 Guide Star Acquisitions:

Scheduled Acquisitions: 9

Successful: 9

Scheduled Re-acquisitions: 6

Successful: 6

2.2 FHST Updates:

Scheduled: 14

Successful: 14

2.3 Operations Notes:

After WF/PC-2 was recovered at 066/0400Z, the instrument began
recording images as 066/073543Z. During the Solid State Recorder dump at
066/212016Z, errors occurred with messages indicating that there were
incomplete packets. Later, at 067/0215Z, PACOR reported that there were
missing headers. Recovery from White Sands was attempted
unsuccessfully. The data is planned to be re-dumped at 067/1020Z. HSTAR
8110 documents these events.

One three occasions early on day 067, telemetry format “P” was
activated for SAA passages.

Using ROP SR-1A, the SSR EDAC error counter was cleared at 067/2305Z.


Continuation of normal science observations and calibrations.

SpaceRef staff editor.